OpenStudy (anonymous):

What is the value of 5 to the power of 4 over 5 to the power of 6? 1 over 25 1 over 5 5 25

OpenStudy (study312):

so\[5^{\frac{ 4 }{ 5^{6} }}\]?

OpenStudy (anonymous):

5^4 over 5^6

OpenStudy (phi):

5^4 is short hand for 5*5*5*5 and 5^6 is short for 5*5*5*5*5*5 can you simplify \[ \frac{5\cdot 5\cdot 5\cdot 5}{5\cdot 5\cdot 5\cdot 5\cdot 5\cdot 5}\]?

OpenStudy (anonymous):

4 over 6

OpenStudy (phi):

do you remember how you multiply fractions? top times top and bottom times bottom for example, we could write that as \[ \frac{5}{5} \cdot \frac{5\cdot 5\cdot 5}{5\cdot 5\cdot 5\cdot 5\cdot 5}=\frac{5\cdot 5\cdot 5\cdot 5}{5\cdot 5\cdot 5\cdot 5\cdot 5\cdot 5} \] (we could write it lots of ways, but notice if you multiply top times top and bottom times bottom you get 5^4/5^6 but notice, what is 5/5 ?

OpenStudy (phi):

do you know that anything divided by itself is 1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait so 1 over 2

OpenStudy (phi):

if we wanted we could "break up" the fraction into lots of 5/5 fractions that multiply together. Each 5/5 turns into 1 **wait so 1 over 2** almost, but it is 1/(5*5) (not 2)

OpenStudy (phi):

If you have a 5 "up top" and a 5 "below" you have a 5/5 and that is 1 so you can "cancel" that pair.

OpenStudy (phi):

If you get the idea that your problem is the *same* as \[ \frac{5}{5} \cdot \frac{5}{5} \cdot \frac{5}{5} \cdot \frac{5}{5} \cdot \frac{1}{5\cdot 5} \] and each 5/5 is 1 can you can simplify that ?

OpenStudy (anonymous):

sorry just got back had to do something

OpenStudy (anonymous):

1 over 5

OpenStudy (phi):

in your problem \[ \frac{5}{5} \cdot \frac{5}{5} \cdot \frac{5}{5} \cdot \frac{5}{5} \cdot \frac{1}{5\cdot 5} \] change each 5/5 into 1 (because 5 divided by itself is 1) now multiply it out what do you get ?